# Sample Efficient Algorithms for Linear System Identification under Noisy Observations

### v1

@@ IV-B Theoretical Guarantee
Now we present the theoretical guarantee, along with necessary assumptions, of the BC estimator.
@@
Consider system $\mathcal{M}=(A,B,\Sigma_{w},\Sigma_{u},\Sigma_{\eta})$ that satisfies Assumption 2 with constant $\phi_{R}$. Suppose $\sigma_{m}\left(\Sigma_{u}\right)\coloneqq\phi_{u}$ satisfy
@@ V Simulations
With a single trajectory from the above system with length $T$, we computed the naive least-squares estimator (Equation 2), the Ho-Kalman estimator [15], the IV estimator (LABEL:eq:iv_estimator) and the BC estimator (Equation 18). As is shown in Figure 1, errors of the proposed IV and BC estimators converge to zero as $T$ while the naive least-squares estimator remains biased. Moreover, the proposed algorithms converge much faster than the Ho-Kalman algorithm in this setting.
The autonomous system shares the same $A,\Sigma_{w},\Sigma_{\eta}$ and $\Sigma_{u}$. With a single trajectory of length $T$ from the system, we computed the naive least-squares estimator (Remark 2), the IV estimator (Remark 1) and the BC estimator (Remark 2). The errors of the proposed IV and BC estimators converge to 00 while the naive least-squares estimator remains biased.
@@ Appendix A Proofs for Section III
$\begin{split}\sum_{t=1}^{T-1}i_{t}i_{t}^{\intercal}={}&\sum_{t=1}^{T-1}\begin{bmatrix}x_{t-1}\\ u_{t}\end{bmatrix}\begin{bmatrix}x_{t-1}\\ u_{t}\end{bmatrix}^{\intercal}=\sum_{t=1}^{T-1}\begin{bmatrix}x_{t-1}x_{t-1}^{\intercal}&0\\ 0&u_{t}u_{t}^{\intercal}\end{bmatrix}+\begin{bmatrix}0&x_{t-1}u_{t}^{\intercal}\\ u_{t}x_{t-1}^{\intercal}&0\end{bmatrix}.\end{split}$ (22)
@@
Substituting Equations 26, 23 and 24 back into Equation 22 gives the following for some absolute constant $c_{1}$ with probability at least $1-4\delta^{\prime}$
@@
$\begin{split}\widehat{z}_{t}={}&\begin{bmatrix}\widehat{x}_{t}\\ u_{t}\end{bmatrix}=\begin{bmatrix}Ax_{t-1}+Bu_{t-1}+w_{t-1}+\eta_{t}\\ u_{t}\end{bmatrix}\\ ={}&\begin{bmatrix}A&0\\ 0&I\end{bmatrix}i_{t}+\begin{bmatrix}Bu_{t-1}+w_{t-1}+\eta_{t}\\ 0\end{bmatrix}.\end{split}$ (40)
@@
$\begin{split}\left\|\sum_{t=1}^{T-1}\eta_{t-1}u_{t}^{\intercal}\right\|,\left\|\sum_{t=1}^{T-1}\eta_{t-1}u_{t-1}^{\intercal}\right\|\lesssim{}&\sqrt{\psi_{u}\psi_{\eta}nT\log\left(\frac{\psi_{\eta}n}{\delta^{\prime}}\right)},\\ \left\|\sum_{t=1}^{T-1}\eta_{t-1}w_{t-1}^{\intercal}\right\|\lesssim{}&\sqrt{\psi_{w}\psi_{\eta}nT\log\left(\frac{\psi_{\eta}n}{\delta^{\prime}}\right)},\\ \left\|\sum_{t=1}^{T-1}u_{t-1}u_{t}^{\intercal}\right\|\lesssim{}&\sqrt{\psi_{u}^{2}nT\log\left(\frac{\psi_{u}n}{\delta^{\prime}}\right)}\\ \left\|\sum_{t=1}^{T-1}u_{t}w_{t-1}^{\intercal}\right\|\lesssim{}&\sqrt{\psi_{u}\psi_{w}nT\log\left(\frac{\psi_{u}n}{\delta^{\prime}}\right)}\\ \end{split}$ (43)
@@
Combining Equations 42, 43, 44 and 45, we have the following with probability at least $1-8\delta^{\prime}$,
@@
Now let $\Sigma_{i}\coloneqq\sum_{t=1}^{T-1}i_{t}i_{t}^{\intercal}$ Substituting back into Equation 41 gives the following with probability at least $1-8\delta^{\prime}$
$\begin{split}{}&\sigma_{m+n}\left(\left(\sum_{t=1}^{T-1}\widehat{z}_{t}\widehat{i}_{t}^{\intercal}\right)\Sigma_{i}^{-1}\right)\\ \geq{}&\sigma_{m+n}\left(\begin{bmatrix}A&0\\ 0&I\end{bmatrix}\right)-\sigma_{1}\left(\sum_{t=1}^{T-1}\begin{bmatrix}Bu_{t-1}+w_{t-1}+\eta_{t}\\ 0\end{bmatrix}\begin{bmatrix}x_{t-1}\\ u_{t}\end{bmatrix}^{\intercal}\Sigma_{i}^{-1}\right)\\ {}&-\sigma_{1}\left(\sum_{t=1}^{T-1}\begin{bmatrix}x_{t-1}\\ u_{t}\end{bmatrix}\begin{bmatrix}\eta_{t-1}\\ 0\end{bmatrix}^{\intercal}\Sigma_{i}^{-1}\right)-\sigma_{1}\left(\sum_{t=1}^{T-1}\begin{bmatrix}Bu_{t-1}+w_{t-1}+\eta_{t}\\ 0\end{bmatrix}\begin{bmatrix}\eta_{t-1}\\ 0\end{bmatrix}^{\intercal}\Sigma_{i}^{-1}\right)\\ \geq{}&\min\{\phi_{A},1\}-\left\|\sum_{t=1}^{T-1}\begin{bmatrix}Bu_{t-1}+w_{t-1}+\eta_{t}\\ 0\end{bmatrix}\begin{bmatrix}x_{t-1}\\ u_{t}\end{bmatrix}^{\intercal}\right\|\left\|\Sigma_{i}^{-1}\right\|\\ {}&-\left\|\sum_{t=1}^{T-1}\begin{bmatrix}x_{t-1}\\ u_{t}\end{bmatrix}\begin{bmatrix}\eta_{t-1}\\ 0\end{bmatrix}^{\intercal}\right\|\left\|\Sigma_{i}^{-1}\right\|-\left\|\sum_{t=1}^{T-1}\begin{bmatrix}Bu_{t-1}+w_{t-1}+\eta_{t}\\ 0\end{bmatrix}\begin{bmatrix}\eta_{t-1}\\ 0\end{bmatrix}^{\intercal}\right\|\left\|\Sigma_{i}^{-1}\right\|\\ \end{split}$ (48)
@@ A-A Supporting Details
[Adapted from Corollary 8.2 in Sasha] Suppose $A$ satisfy Assumption 1 with constant $\psi_{A},\rho_{A}$. Then the following holds with probability at least $1-\delta$ for any $\delta\in(0,e^{-1})$
@@
Directly applying Proposition 8.4 in Sasha gives the following with probability at least $1-\delta$
@@ Appendix B Proof of Theorem 2
$\begin{split}{}&\widehat{E}_{\text{BC}}-E=\widehat{E}_{\text{LS}}\left(I-T\widehat{\Sigma}_{\tilde{\eta}}\left(\Sigma_{\widehat{z}}\right)^{-1}\right)^{-1}-E\\ ={}&E\left(I-\left(\sum_{t=0}^{T-1}\tilde{\eta}_{t}\tilde{\eta}_{t}^{\intercal}\right)\left(\Sigma_{\widehat{z}}\right)^{-1}\right)\left(I-T\widehat{\Sigma}_{\tilde{\eta}}\left(\Sigma_{\widehat{z}}\right)^{-1}\right)^{-1}-E+\Delta_{1}\left(I-T\widehat{\Sigma}_{\tilde{\eta}}\left(\Sigma_{\widehat{z}}\right)^{-1}\right)^{-1}\\ ={}&E\left(T\widehat{\Sigma}_{\tilde{\eta}}-\sum_{t=0}^{T-1}\tilde{\eta}_{t}\tilde{\eta}_{t}^{\intercal}\right)\left(\Sigma_{\widehat{z}}\right)^{-1}\left(I-T\widehat{\Sigma}_{\tilde{\eta}}\left(\Sigma_{\widehat{z}}\right)^{-1}\right)^{-1}+\Delta_{1}\left(I-T\widehat{\Sigma}_{\tilde{\eta}}\left(\Sigma_{\widehat{z}}\right)^{-1}\right)^{-1}\\ \end{split}$ (58)
@@
Substituting Equations 56, 57 and 59 into LABEL:eq:errdecomp_bc gives the following with probability at least $1-3\delta$
@@
Therefore, combining Equations 61 and 64, we have the following with probability at least $1-3\delta$
@@
Finally, combining LABEL:eq:bcmain_6, 62 and 57 gives the following with probability at least $1-6\delta$
@@ B-A Supporting Details
$\begin{split}\sum_{t=0}^{T-1}z_{t}z_{t}^{\intercal}={}&\sum_{t=0}^{T-1}\begin{bmatrix}x_{t}\\ u_{t}\end{bmatrix}\begin{bmatrix}x_{t}\\ u_{t}\end{bmatrix}^{\intercal}=\sum_{t=0}^{T-1}\begin{bmatrix}x_{t}x_{t}^{\intercal}&0\\ 0&u_{t}u_{t}^{\intercal}\end{bmatrix}+\begin{bmatrix}0&x_{t}u_{t}^{\intercal}\\ u_{t}x_{t}^{\intercal}&0\end{bmatrix}.\end{split}$ (68)
@@
Substituting Equations 72, 69 and 70 back into Equation 68 gives the following for some absolute constant $c_{1}$ with probability at least $1-4\delta^{\prime}$
@@
$\begin{split}\sum_{t=0}^{T-1}\widehat{z}_{t}\widehat{z}_{t}^{\intercal}={}&\sum_{t=0}^{T-1}z_{t}z_{t}^{\intercal}+z_{t}\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}^{\intercal}+\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}z_{t}^{\intercal}+\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}^{\intercal}.\end{split}$ (77)
@@
$\begin{split}{}&\sigma_{m+n}\left(\left(\sum_{t=0}^{T-1}\widehat{z}_{t}\widehat{z}_{t}^{\intercal}\right)\overline{\Sigma}_{z}^{-1}\right)\\ \geq{}&1-\sigma_{1}\left(\left(\sum_{t=0}^{T-1}z_{t}\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}^{\intercal}+\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}z_{t}^{\intercal}\right)\overline{\Sigma}_{z}^{-1}\right)-\sigma_{1}\left(\left(\sum_{t=0}^{T-1}\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}\begin{bmatrix}\eta_{t}\\ 0\end{bmatrix}^{\intercal}\right)\overline{\Sigma}_{z}^{-1}\right)\\ \geq{}&1-\frac{15\psi_{\eta}}{\min\left\{\phi_{R}^{2},6\right\}\phi_{u}}\\ {}&-c_{1}\sqrt{\frac{\psi^{2}\psi_{A}^{2}}{(1-\rho_{A}^{2})\phi_{u}^{2}\min\left\{\phi_{R}^{4},1\right\}}\log\left(\frac{1}{\delta^{\prime}}\right)\log\left(\frac{\psi\psi_{A}^{2}n}{\left(1-\rho_{A}^{2}\right)\delta^{\prime}}\log\frac{1}{\delta^{\prime}}\right)}\sqrt{\frac{(m+n)^{2}}{T}}\\ \end{split}$ (86)

### v2

@@ IV-B Theoretical Guarantee
Now we present the error upper bound for the BC estimator (Equation 19), along with necessary assumptions.
@@
Consider system $\mathcal{M}=(A,B,\Sigma_{w},\Sigma_{u},\Sigma_{\eta})$ that satisfies Assumption 2 with constant $\phi_{R}$. Suppose $\sigma_{m}\left(\Sigma_{u}\right)\coloneqq\phi_{u}$ satisfies
@@ V Simulations
With a single trajectory from the above system with length $T$, we computed the naive least-squares estimator (Equation 2), the Ho-Kalman estimator [15], the IV estimator (LABEL:eq:iv_estimator) and the BC estimator (Equation 19). As is shown in Figure 1, errors of the proposed IV and BC estimators converge to almost zero as $T$ while the naive least-squares estimator remains biased. Moreover, the proposed algorithms converge much faster than the Ho-Kalman algorithm in this setting.
The autonomous system shares the same $A,\Sigma_{w},\Sigma_{\eta}$ and $\Sigma_{u}$. With a single trajectory of length $T$ from the system, we computed the naive least-squares estimator (Remark 2), the IV estimator (Remark 1) and the BC estimator (Remark 2). The errors of the proposed IV and BC estimators converge to almost 0 while the naive least-squares estimator remains biased.
@@ Appendix A Proofs for Section III
[…]
@@
Substituting Equations 27, 24 and 25 back into Equation 23 gives the following for some absolute constant $c_{1}$ with probability at least $1-4\delta^{\prime}$
@@
[…]
@@
[…]
@@
Combining Equations 43, 44, 45 and 46, we have the following with probability at least $1-8\delta^{\prime}$,
@@
Now let $\Sigma_{i}\coloneqq\sum_{t=1}^{T-1}i_{t}i_{t}^{\intercal}$ . Substituting back into Equation 42 gives the following with probability at least $1-8\delta^{\prime}$
[…]
@@ A-A Supporting Details
[Adapted from Corollary 8.2 in [21]] Suppose $A$ satisfies Assumption 1 with constant $\psi_{A},\rho_{A}$. Then the following holds with probability at least $1-\delta$ for any $\delta\in(0,e^{-1})$
@@
Directly applying Proposition 8.4 in [21] gives the following with probability at least $1-\delta$
@@ Appendix B Proof of Theorem 2
[…]
@@
Substituting Equations 57, 58 and 60 into LABEL:eq:errdecomp_bc gives the following with probability at least $1-3\delta$
@@
Therefore, combining Equations 62 and 65, we have the following with probability at least $1-3\delta$
@@
Finally, combining LABEL:eq:bcmain_6, 63 and 58 gives the following with probability at least $1-6\delta$
@@ B-A Supporting Details
[…]
@@
Substituting Equations 73, 70 and 71 back into Equation 69 gives the following for some absolute constant $c_{1}$ with probability at least $1-4\delta^{\prime}$
@@
[…]
@@
[…]